# Orbit-based structural decomposition and stellar population recovery for edge-on barred galaxies

### v1

@@ 3 Barred population-orbit superposition models
Figure 2: Structural decomposition based on orbital properties demonstrated using Au-23-85-50 and compared with the truth. Four properties are used to decompose the galaxy: circularity $\lambda_{z}$, time-averaged radius $R$, axis ratio $p_{\rm orb}$ in the $x$-$y$ plane, and axis ratio $q_{\rm orb}$ in the $x$-$z$ plane. The top panels display the model results while the bottom panels correspond to the truth. Left panels: the stellar orbit distributions of the entire galaxy in the $\lambda_{z}$$R$ phase space. We calculate the 1-kpc moving average of the cold orbit fraction $f_{\rm cold}$ ($\lambda_{z}\geq 0.8$; $\rm 1\,arcsec=0.2\,kpc$), and define the dynamical bar length $R_{\rm bar}$ as the smallest radius where $f_{\rm cold}\geq 0.5$. Orbits with $R\leq R_{\rm bar}$ are classified as bar/bulge components; those with $R>R_{\rm bar}$ are categorized as disk/halo components. The values of $R_{\rm bar}$ are indicated by the blue dashed lines and the annotations. Middle panels: the stellar orbit distributions for bar and bulge components in the $\lambda_{z}$$p_{\rm orb}$ phase space. Orbits with $\lambda_{z}\leq 0.5$ and $p_{\rm orb}\geq 0.8$ are assigned to the bulge; other orbits constitute the bar. The boundaries between the bar and bulge are shown by the blue dashed lines. Right panels: the stellar orbit distributions for disk and halo components in the $\lambda_{z}$$q_{\rm orb,in}$ phase space. Orbits with $\lambda_{z}\leq 0.5$ and $q_{\rm orb,in}\geq 0.3$ are categorised as the stellar halo; other orbits make up the disk. The boundaries between the disk and halo are shown by the blue dashed lines.
@@
Figure 3: The trajectories of several representative orbits for Au-23-85-50. The first four columns display the orbital trajectories in the $x$-$y$, $x$-$z$, and $y$-$z$ planes (the bar’s rotating frame; $\rm 1\,arcsec=0.2\,kpc$), including two bar orbits and two bulge orbits, with their axis ratios $p_{\rm orb}$ and $q_{\rm orb}$ presented in the text. The fifth and sixth columns show the orbital trajectories in the $x_{\rm in}$-$y_{\rm in}$, $x_{\rm in}$-$z_{\rm in}$, and $y_{\rm in}$-$z_{\rm in}$ planes (the inertial frame), including a disk orbit and a halo orbit, with their axis ratios $p_{\rm orb,in}$ and $q_{\rm orb,in}$ presented in the text. We note that $p_{\rm orb}$ and $q_{\rm orb}$ are calculated in the bar’s rotating frame, while $\lambda_{z}$, $p_{\rm orb,in}$, and $q_{\rm orb,in}$ are calculated in the inertial frame.
@@
Figure 4: True and model-predicted probability density distributions of stellar orbits in the $\lambda_{z}$$R$ phase space for different structures in Au-23. The top panels indicate the true distributions while the bottom panels represent the model results from Au-23-85-50. From left to right, the panels show the stellar orbit distributions for the bar, the bulge, the disk, and the stellar halo. The probability densities of all orbits within the data coverage ($R\leq\rm 50\,arcsec$; $\rm 1\,arcsec=0.2\,kpc$) are normalised to unity, with their values indicated by the colour bar. The horizontal black dotted lines denote $\lambda_{z}=0$, and the vertical blue dashed lines represent the true and model-predicted bar lengths $R_{\rm bar}$ derived from orbit analysis, with their values shown in the text.
@@
Figure 5: The orbit bundles divided in the phase space of circularity $\lambda_{z}$ versus time-averaged radius $R$ for Au-23-85-50. The left panel represents bar orbits while the right panel is for orbits from other components, including the bulge, disk, and stellar halo. The probability densities of all orbits within the data coverage ($R\leq\rm 50\,arcsec$; $\rm 1\,arcsec=0.2\,kpc$) are normalised to unity, with their values indicated by the colour bar. For each panel, orbits in $\lambda_{z}$$R$ phase space are divided into different bundles by using the Voronoi binning method, with each bundle containing $\gtrsim 0.5\%$ of the total orbit weight. The blue asterisks and lines indicate the centres and boundaries of Voronoi bins, respectively.
@@
Figure 6: Mock surface brightness, stellar kinematic, age, and metallicity maps (left panels), best-fitting model for Au-23-85-50 (middle panels), and residuals (right panels). From top to bottom: Normalised surface brightness $\log\Sigma$, mean velocity $V$, velocity dispersion $\sigma$, third-order $h_{3}$ and fourth-order $h_{4}$ Gauss-Hermite coefficients, stellar ages $t$, and stellar metallicity $Z$ ($\rm 1\,arcsec=0.2\,kpc$). Residuals for surface brightness are relative deviations, while those for kinematics and stellar populations are standardised residuals. The surface brightness and kinematic maps are the same as Fig. 7 of Jin et al. (2025).
@@
We propose a three-step methodology to construct barred population-orbit superposition models. First, we create barred orbit-superposition models by simultaneously fitting the stellar luminosity distributions and kinematic data, and search for the best-fitting model (Sect. 3.1). This step builds directly on our previous work (Jin et al., 2025), which contains detailed explanations, and we only briefly summarise it here. Then, we categorise the stellar orbits of the best-fitting model into different components based on their orbital properties (Sect. 3.2). Finally, we tag the orbits with ages and metallicities by fitting the stellar age and metallicity maps (Sect. 3.3).
@@ 3.1 Barred orbit-superposition models
The procedures for creating a barred orbit-superposition model include three major steps: (1) constructing the gravitational potential using a set of free parameters, (2) sampling and integrating orbits, and (3) solving orbit weights by fitting the observational data. By exploring the free parameter space, the above steps are repeated until the best-fitting model is found.
@@
The gravitational potential of a barred galaxy contains three components: (1) a triaxial stellar bar with constant figure rotation $\rm\Omega_{p}$, (2) an axisymmetric stellar disk with its major axis aligned with the bar, and (3) a spherical dark matter halo. The central black hole is not considered in the modelling due to the limited spatial resolution ($\rm\sim 0.2\,kpc/pixel$) of the mock data. To calculate the stellar potential, the galaxy’s surface brightness is first fitted by the multi-Gaussian expansion (MGE) formalism (Emsellem et al., 1994; Cappellari, 2002), with each Gaussian being assigned either to the triaxial bar component or to the axisymmetric disk component. The 2D Gaussian components of the bar and disk are separately deprojected into the 3D luminosity density using distinct viewing angles: the inclination angle $\theta_{\rm disk}$ and the position angle $\psi_{\rm disk}$ for the axisymmetric disk and the inclination angle $\theta_{\rm bar}$, the bar azimuthal angle $\varphi$, and the position angle $\psi_{\rm bar}$ for the triaxial bar. Taking the assumption that the bar major axis ($x$-axis) is in the disk plane, the inclination angles of both components become equal ($\theta_{\rm bar}=\theta_{\rm disk}=\theta$). The disk in the observing plane is rotated to make sure its major axis aligns with the $x^{\prime}$-axis ($\psi_{\rm disk}=90^{\circ}$). For edge-on views, the bar position angle $\psi_{\rm bar}$, which represents the misalignment between the bar major axis and the disk major axis in the observing plane, is difficult to measure directly. However, small changes in $\psi_{\rm bar}$ will significantly alter the morphology of the deprojected bar. Therefore, $\psi_{\rm bar}$ is determined in each set of ($\theta,\varphi$) by adding physical constraints to the bar morphology, and thus only the viewing angles $\theta$ and $\varphi$ remain free during the deprojection (see Sect. 3.1.1 in Jin et al., 2025 for details). By multiplying the 3D mass density by a constant mass-to-light ratio $M_{*}/L$ and solving Poisson’s equation, both the potentials of the stellar bar and the stellar disk are derived. We note that the bar and the disk here are used to construct the stellar potential and are distinct from the dynamically decomposed bar and disk mentioned later in Sect. 3.2. The spherical dark matter potential is described by a three-parameter generalised NFW (gNFW) profile (Navarro et al., 1996; Zhao, 1996; see also Barnabè et al., 2012; Cappellari et al., 2013). Overall, there are seven free parameters in the modelling: the inclination angle $\theta$, the bar azimuthal angle $\varphi$, the stellar mass-to-light ratio $M_{*}/L$, the bar pattern speed $\rm\Omega_{p}$, the dark matter concentration $c$, the virial mass $M_{200}$, and the inner density slope of dark matter $\gamma$.
@@
The orbit weights are solved using the non-negative least squares (NNLS; Lawson & Hanson, 1974) method, taking the 2D surface brightness fitted by MGE, the 3D mass density deprojected from MGE, and the kinematic data as constraints. After solving the orbit weights, the goodness of the model is evaluated by calculating the difference between the model-fitted and observed kinematic maps, which is expressed as
@@ 3.2 Decomposing galaxies based on stellar orbits
The best-fitting model contains an ensemble of stellar orbits. In this paper, orbits are categorised into four distinct components: bar, bulge, disk, and stellar halo, based on four orbital properties, including (1) circularity $\lambda_{z}$, (2) time-averaged radius $R$, (3) axis ratio $p_{\rm orb}$ in the $x$-$y$ plane (bar’s rotating frame), and (4) axis ratio $q_{\rm orb,in}$ in the $x_{\rm in}$-$z_{\rm in}$ plane (inertial frame). We note that circularity is defined exclusively in the inertial frame ($x_{\rm in},y_{\rm in},z_{\rm in}$), while the axis ratios are calculated in both the inertial frame and the bar’s rotating frame ($x,y,z$).
The circularity represents the orbit’s angular momentum, which is defined as a ratio of time-averaged quantities
$\lambda_{z}=\overline{L_{z}}/(\overline{R}\times\overline{V_{\rm rms}}),$ (2)
where $\overline{L_{z}}=\overline{x_{\rm in}v_{y\rm in}-y_{\rm in}v_{x\rm in}}$, $\overline{R}=\overline{\sqrt{x_{\rm in}^{2}+y_{\rm in}^{2}+z_{\rm in}^{2}}}$ and $\overline{V_{\rm rms}}=\sqrt{\overline{v_{x\rm in}^{2}+v_{y\rm in}^{2}+v_{z\rm in}^{2}+2v_{x\rm in}v_{y\rm in}+2v_{x\rm in}v_{z\rm in}+2v_{y\rm in}v_{z\rm in}}}$. Orbits with $\lambda_{z}\sim 1$ are rotation-dominated, orbits with $\lambda_{z}\sim 0$ are dispersion-dominated, and orbits with $\lambda_{z}<1$ correspond to retrograde orbits.
Orbit circularity $\lambda_{z}$ and radius $R$ have been widely used in previous studies to decompose galaxies without considering the bars (e.g. Zhu et al., 2018a, b, c, 2020, 2022a, 2022b, 2025; Jin et al., 2019, 2020, 2024; Poci et al., 2019, 2021; Santucci et al., 2022, 2023, 2024; Ding et al., 2023, 2024; Zhang et al., 2025). However, for barred galaxies, these criteria alone fail to separate the bulge and bar in the innermost regions, as both structures exhibit orbits with $\lambda_{z}\sim 0$. Additionally, in the outer regions of galaxies, there are non-circular disk orbits with obvious radial motion ($\lambda_{z}\ll 1$), which overlap the circularity range of halo orbits.
To break the degeneracy between different structures in the $\lambda_{z}$$R$ phase space, we further introduce the axis ratios $p_{\rm orb}=b/a$ and $q_{\rm orb}=c/a$ calculated from orbital trajectories, with $a$, $b$, and $c$ denoting the time-averaged amplitudes along the $x$-axis, $y$-axis, and $z$-axis, respectively. Thus, $p_{\rm orb}$ and $q_{\rm orb}$ are expressed as
@@
Similarly, we calculate the axis ratios $p_{\rm orb,in}$ and $q_{\rm orb,in}$ of orbits in the inertial frame ($x_{\rm in},y_{\rm in},z_{\rm in}$).
@@
From the kinematic perspective, the disk is dominated by dynamically cold orbits ($\lambda_{z}\geq 0.8$), whereas the bar rotates as a rigid body with low $\lambda_{z}$ in its inner region and becomes corotating with the disk at the bar end. Therefore, we define the dynamical bar length $R_{\rm bar}$ as the radius where cold orbits begin to dominate. We calculate the 1-kpc moving average of the cold orbit fraction $f_{\rm cold}$, and identify $R_{\rm bar}$ as the smallest radius where $f_{\rm cold}\geq 0.5$. In this way, we obtain $R_{\rm bar,model}=\rm 23.1\,arcsec$ ($\rm 1\,arcsec=0.2\,kpc$) for the model of Au-23-85-50.
@@
For each simulated galaxy, we freeze all the stellar and dark matter particles in the redshift $z=0$ snapshot and calculate the gravitational potential from their distributions using multipole expansion of spherical harmonics (Binney & Tremaine, 2008). Taking the 6D phase-space information of stellar particles as initial conditions, we integrate orbits for $\sim 100$ orbital periods in the bar’s rotating frame (with true pattern speeds from Fragkoudi et al., 2020, 2021) via the AGAMA software (Vasiliev, 2019) [^2] https://github.com/GalacticDynamics-Oxford/Agama, and calculate the same parameters characterising the orbits. The true bar length $R_{\rm bar}$ is then determined the same way as in the model, and we obtain $R_{\rm bar,true}=\rm 24.6\,arcsec$ for Au-23.
@@
We identify different structures based on the orbital information using identical methodology in both our models and the simulations. We first divide the orbits with $R\leq R_{\rm bar}$ as bar/bulge components and those with $R>R_{\rm bar}$ as disk/halo components, using $R_{\rm bar,model}$ and $R_{\rm bar,true}$ for our models and the simulations, respectively. Then we further classify different components according to the following criteria: (1) bar orbits are rotation-dominated ($\lambda_{z}>0.5$) and/or elongated along the bar ($p_{\rm orb}<0.8$), (2) bulge orbits are random-motion-dominated ($\lambda_{z}\leq 0.5$), and not elongated along the bar ($p_{\rm orb}\geq 0.8$), (3) disk orbits are rotation-dominated ($\lambda_{z}>0.5$) and/or thin ($q_{\rm orb,in}<0.3$), and (4) halo orbits are random-motion-dominated ($\lambda_{z}\leq 0.5$) and thick ($q_{\rm orb,in}\geq 0.3$). In Fig. 2, we demonstrate these criteria using Au-23-85-50. The trajectories of several representative orbits belonging to different structures for Au-23-85-50 are illustrated in Fig. 3.
@@
We show the probability density distributions of stellar orbits for Au-23 and those recovered by the model of Au-23-85-50 in Fig. 4. We successfully decompose the galaxy into a rigidly rotating bar ($\lambda_{z}$ is strongly correlated with $R$), a random-motion-dominated bulge ($|\lambda_{z}|\lesssim 0.5$), a rotation-dominated disk ($\lambda_{z}\gtrsim 0.8$), and a stellar halo; the stellar orbit distributions of these components are similar in the true simulation and our model.

### v4

@@ 3 Barred population-orbit superposition models
Figure 2: Structural decomposition based on orbital properties demonstrated using Au-23-85-50 and compared with the truth. Four properties are used to decompose the galaxy: Circularity, $\lambda_{z}$; time-averaged radius, $R$; axis ratio, $p_{\rm orb}$, in the $x$-$y$ plane; and axis ratio, $q_{\rm orb}$, in the $x$-$z$ plane. The top panels display the true distributions while the bottom panels correspond to the model results. Left panels: Stellar orbit distributions of the entire galaxy in the $\lambda_{z}$$R$ phase space. We calculate the $\rm 1\,kpc$ moving average of the cold orbit fraction $f_{\rm cold}$ ($\lambda_{z}\geq 0.8$; $\rm 1\,arcsec=0.2\,kpc$) and define the dynamical bar length, $R_{\rm bar}$, as the smallest radius where $f_{\rm cold}\geq 0.5$. Orbits with $R\leq R_{\rm bar}$ are classified as bar+bulge components; those with $R>R_{\rm bar}$ are categorised as disc+halo components. The values of $R_{\rm bar}$ are indicated by the blue dashed lines and the annotations. Middle panels: Stellar orbit distributions for bar and bulge components in the $\lambda_{z}$$p_{\rm orb}$ phase space. Orbits with $\lambda_{z}\leq 0.5$ and $p_{\rm orb}\geq 0.8$ are assigned to the bulge; other orbits constitute the bar. The boundaries between the bar and bulge are shown by the blue dashed lines. Right panels: Stellar orbit distributions for disc+halo components in the $\lambda_{z}$$q_{\rm orb,in}$ phase space. Orbits with $\lambda_{z}\leq 0.5$ and $q_{\rm orb,in}\geq 0.3$ are categorised as the stellar halo; while the other orbits make up the disc. The boundaries between the disc and halo are shown by the blue dashed lines.
@@
Figure 3: Trajectories of several representative orbits for Au-23-85-50. The first four columns display the orbital trajectories in the $x$-$y$, $x$-$z$, and $y$-$z$ planes (the bar’s rotating frame), including two bar orbits and two bulge orbits, with their axis ratios $p_{\rm orb}$ and $q_{\rm orb}$ presented in the text. The fifth and sixth columns show the orbital trajectories in the $x_{\rm in}$-$y_{\rm in}$, $x_{\rm in}$-$z_{\rm in}$, and $y_{\rm in}$-$z_{\rm in}$ planes (the inertial frame), including a disc orbit and a halo orbit, with their axis ratios, $p_{\rm orb,in}$ and $q_{\rm orb,in}$, presented in the text. We note that $p_{\rm orb}$ and $q_{\rm orb}$ are calculated in the bar’s rotating frame, while $\lambda_{z}$, $p_{\rm orb,in}$, and $q_{\rm orb,in}$ are calculated in the inertial frame.
@@
Figure 4: True and model-predicted probability density distributions of stellar orbits in the $\lambda_{z}$$R$ phase space for different structures in Au-23. The top panels indicate the true distributions while the bottom panels represent the model results from Au-23-85-50. From left to right, the panels show the stellar orbit distributions for the bar, the bulge, the disc, and the stellar halo. The probability densities of all orbits within the data coverage ($R\leq\rm 50\,arcsec$) are normalised to unity, with their values indicated by the colour bar. The horizontal black dotted lines denote $\lambda_{z}=0$. The vertical blue dashed lines represent the true and model-predicted bar lengths, $R_{\rm bar}$, derived from orbit analysis, with their values shown in the text.
@@
Figure 5: Orbit bundles divided in the phase space of circularity, $\lambda_{z}$, versus time-averaged radius, $R$, for Au-23-85-50. The left panel represents bar orbits while the right panel is for orbits from other components, including the bulge, disc, and stellar halo. The probability densities of all orbits within the data coverage ($R\leq\rm 50\,arcsec$) are normalised to unity, with their values indicated by the colour bar. For each panel, orbits in $\lambda_{z}$$R$ phase space are divided into different bundles by using the Voronoi binning method, with each bundle containing $\gtrsim 0.5\%$ of the total orbit weight. The blue asterisks and lines indicate the centres and boundaries of Voronoi bins, respectively.
@@
Figure 6: Mock surface brightness, stellar kinematic, age, and metallicity maps (left panels), best-fitting model for Au-23-85-50 (middle panels), and residuals (right panels). From top to bottom: Normalised surface brightness, $\log\Sigma$; mean velocity, $V$; velocity dispersion, $\sigma$; third-order, $h_{3}$, and fourth-order, $h_{4}$, Gauss-Hermite coefficients; stellar age, $t$; and stellar metallicity, $Z$. The residuals for surface brightness are relative deviations, while those for kinematics and stellar populations are standardised residuals. The surface brightness and kinematic maps are the same as Fig. 6 of Jin et al. (2025).
@@
We propose a three-step methodology to construct barred population-orbit superposition models. First, we create barred orbit-superposition models by simultaneously fitting the stellar luminosity distributions and kinematic data, and search for the best-fitting model (Sect. 3.1). This step builds directly on our previous work (Jin et al., 2025), which contains detailed explanations and so, we only briefly summarise it here. Then, we categorise the stellar orbits of the best-fitting model into different components based on their orbital properties (Sect. 3.2). Finally, we tag the orbits with ages and metallicities by fitting the stellar age and metallicity maps (Sect. 3.3).
@@ 3.1 Barred orbit-superposition models
The procedures for creating a barred orbit-superposition model include three major steps: (1) constructing the gravitational potential using a set of free parameters; (2) sampling and integrating orbits; and (3) solving orbit weights by fitting the observational data. By exploring the free parameter space, the above steps are repeated until the best-fitting model is found.
@@
The gravitational potential of a barred galaxy contains three components: (1) a triaxial stellar bar with constant figure rotation, $\rm\Omega_{p}$; (2) an axisymmetric stellar disc with its major axis aligned with the bar; and (3) a spherical dark matter halo. The central black hole is not considered in the modelling due to the limited spatial resolution ($\rm\sim 0.2\,kpc/pixel$) of the mock data. To calculate the stellar potential, the galaxy’s surface brightness is first fitted by the multi-Gaussian expansion (MGE) formalism (Emsellem et al., 1994; Cappellari, 2002), with each Gaussian being assigned either to the triaxial bar component or to the axisymmetric disc component. The 2D Gaussian components of the bar and disc are separately deprojected into the 3D luminosity density using distinct viewing angles: the inclination angle, $\theta_{\rm disc}$, and the position angle, $\psi_{\rm disc}$, for the axisymmetric disc and the inclination angle, $\theta_{\rm bar}$, the bar azimuthal angle, $\varphi$, and the position angle, $\psi_{\rm bar}$, for the triaxial bar. Taking the assumption that the bar major axis ($x$-axis) is in the disc plane, the inclination angles of both components become equal ($\theta_{\rm bar}=\theta_{\rm disc}=\theta$). The disc in the observing plane is rotated to make sure its major axis aligns with the $x^{\prime}$-axis ($\psi_{\rm disc}=90^{\circ}$). For edge-on views, the bar position angle, $\psi_{\rm bar}$, which represents the misalignment between the bar major axis and the disc major axis in the observing plane, is difficult to measure directly. However, small changes in $\psi_{\rm bar}$ will significantly alter the morphology of the deprojected bar. Therefore, $\psi_{\rm bar}$ is determined in each set of ($\theta,\varphi$) by adding physical constraints to the bar morphology; thus, only the viewing angles $\theta$ and $\varphi$ remain free during the deprojection (see Sect. 3.1.1 in Jin et al., 2025 for details). By multiplying the 3D luminosity density by a constant mass-to-light ratio, $M_{*}/L$, and solving Poisson’s equation, both the potentials of the stellar bar and the stellar disc can be derived. We note that the bar and disc used here to construct the stellar potential are distinct from the dynamically decomposed bar and disc mentioned later (see Sect. 3.2). The spherical dark matter potential is described by a three-parameter generalised NFW (gNFW) profile (Navarro et al., 1996; Zhao, 1996; see also Barnabè et al., 2012; Cappellari et al., 2013). Overall, there are seven free parameters in the modelling: the inclination angle, $\theta$; the bar azimuthal angle, $\varphi$; the stellar mass-to-light ratio, $M_{*}/L$; the bar pattern speed, $\rm\Omega_{p}$; the dark matter concentration, $c$; the virial mass, $M_{200}$; and the inner density slope of dark matter, $\gamma$.
@@
The orbit weights are solved using the non-negative least squares (NNLS; Lawson & Hanson, 1974) method, taking the 2D surface brightness fitted by MGE, the 3D luminosity density deprojected from MGE, and the kinematic data as constraints. After solving the orbit weights, the goodness of the model is evaluated by calculating the difference between the model-fitted and observed kinematic maps, which is expressed as
@@ 3.2 Decomposing galaxies based on stellar orbits
The best-fitting model contains an ensemble of stellar orbits. In this paper, orbits are categorised into four distinct components: bar, bulge, disc, and stellar halo, based on four orbital properties, including (1) circularity, $\lambda_{z}$; (2) time-averaged radius, $R$; (3) axis ratio, $p_{\rm orb}$, in the $x$-$y$ plane (bar’s rotating frame); and (4) axis ratio, $q_{\rm orb,in}$, in the $x_{\rm in}$-$z_{\rm in}$ plane (inertial frame). We note that circularity is defined exclusively in the inertial frame ($x_{\rm in},y_{\rm in},z_{\rm in}$), while the axis ratios are calculated in both the inertial frame and the bar’s rotating frame ($x,y,z$).
The circularity represents the orbit’s angular momentum, which is defined as a ratio of time-averaged quantities,
$\lambda_{z}=\overline{L_{z}}/(R\times\overline{V_{\rm rms}}),$ (2)
where $\overline{L_{z}}=\overline{x_{\rm in}v_{y\rm in}-y_{\rm in}v_{x\rm in}}$, $R=\overline{\sqrt{x_{\rm in}^{2}+y_{\rm in}^{2}+z_{\rm in}^{2}}}$ and $\overline{V_{\rm rms}}=\sqrt{\overline{v_{x\rm in}^{2}+v_{y\rm in}^{2}+v_{z\rm in}^{2}+2v_{x\rm in}v_{y\rm in}+2v_{x\rm in}v_{z\rm in}+2v_{y\rm in}v_{z\rm in}}}$. Orbits with $\lambda_{z}\sim 1$ are rotation-dominated, orbits with $\lambda_{z}\sim 0$ are dispersion-dominated, and orbits with $\lambda_{z}<1$ correspond to retrograde orbits.
Orbit circularity, $\lambda_{z}$, and radius, $R,$ have been widely used in previous studies to decompose galaxies without considering the bars (e.g. Zhu et al., 2018a, b, c, 2020, 2022a, 2022b, 2025; Jin et al., 2019, 2020, 2024; Poci et al., 2019, 2021; Santucci et al., 2022, 2023, 2024; Ding et al., 2023, 2024; Zhang et al., 2025). However, for barred galaxies, these criteria alone fail to separate the bulge and bar in the innermost regions, as both structures exhibit orbits with $\lambda_{z}\sim 0$. Additionally, in the outer regions of galaxies, there are non-circular disc orbits with obvious radial motion ($\lambda_{z}\ll 1$), which overlap with the circularity range of halo orbits.
To break the degeneracy between different structures in the $\lambda_{z}$$R$ phase space, we further introduce the axis ratios, $p_{\rm orb}=b/a$ and $q_{\rm orb}=c/a$, calculated from orbital trajectories, with $a$, $b$, and $c$ denoting the time-averaged amplitudes along the $x$-axis, $y$-axis, and $z$-axis, respectively. Thus, $p_{\rm orb}$ and $q_{\rm orb}$ are expressed as
@@
Similarly, we calculated the axis ratios, $p_{\rm orb,in}$ and $q_{\rm orb,in}$, of the orbits in the inertial frame ($x_{\rm in},y_{\rm in},z_{\rm in}$).
@@
From the kinematic perspective, the disc is dominated by dynamically cold orbits ($\lambda_{z}\geq 0.8$), whereas the bar rotates as a rigid body with low $\lambda_{z}$ in its inner region and becomes corotating with the disc at the bar end. Therefore, we defined the dynamical bar length, $R_{\rm bar}$, as the radius where cold orbits begin to dominate. We calculated the $\rm 1\,kpc$ moving average of the cold orbit fraction, $f_{\rm cold}$, and identified $R_{\rm bar}$ as the smallest radius where $f_{\rm cold}\geq 0.5$. In this way, we obtained $R_{\rm bar,model}=\rm 23.1\,arcsec$ for the model of Au-23-85-50.
@@
For each simulated galaxy, we froze all the stellar and dark matter particles in the redshift $z=0$ snapshot and calculated the gravitational potential from their distributions using multipole expansion of spherical harmonics (Binney & Tremaine, 2008). Taking the 6D phase-space information of stellar particles as the initial conditions, we integrated orbits for $\sim 100$ orbital periods in the bar’s rotating frame (with true pattern speeds from Fragkoudi et al., 2020, 2021) via the AGAMA software (Vasiliev, 2019) [^2] https://github.com/GalacticDynamics-Oxford/Agama. We then calculated the same parameters characterising the orbits. The true bar length, $R_{\rm bar}$, was then determined the same way as in the model and we obtained $R_{\rm bar,true}=\rm 24.6\,arcsec$ for Au-23.
@@
We identified different structures based on orbital information using an identical methodology in both our models and the simulations. We first divided the orbits with $R\leq R_{\rm bar}$ as the bar+bulge components and those with $R>R_{\rm bar}$ as the disc+halo components, using $R_{\rm bar,model}$ and $R_{\rm bar,true}$ for our models and the simulations, respectively. Then we further classified the different components according to the following criteria: (1) bar orbits are rotation-dominated ($\lambda_{z}>0.5$) and/or elongated along the bar ($p_{\rm orb}<0.8$); (2) bulge orbits are random-motion-dominated ($\lambda_{z}\leq 0.5$) and not elongated along the bar ($p_{\rm orb}\geq 0.8$); (3) disc orbits are rotation-dominated ($\lambda_{z}>0.5$) and/or thin ($q_{\rm orb,in}<0.3$); and (4) halo orbits are random-motion-dominated ($\lambda_{z}\leq 0.5$) and thick ($q_{\rm orb,in}\geq 0.3$). In Fig. 2, we exemplify these criteria using Au-23-85-50. The trajectories of several representative orbits belonging to different structures for Au-23-85-50 are illustrated in Fig. 3.
@@
We show the probability density distributions of stellar orbits for Au-23 and those recovered by the model of Au-23-85-50 in Fig. 4. We successfully decomposed the galaxy into a rigidly rotating bar ($\lambda_{z}$ is strongly correlated with $R$), a random-motion-dominated bulge ($|\lambda_{z}|\lesssim 0.5$), a rotation-dominated disc ($\lambda_{z}\gtrsim 0.8$), and a stellar halo; the stellar orbit distributions of these components are similar in the true simulation and our model.